OpenStudy (anonymous):

Calculus Help?????? Sketch the graph of a function f having the given characteristics. f(0)=f(2)=0 f'(x)>0 if x<1 f'(1)=0 f'(x)<0 if x>1 f''(x)<0 I don't know how to do this. Please help??

OpenStudy (anonymous):

A function similar to your function can be defined \[ f(x) = 2 x - x^2\\ \] You cam verify that it would satisfy all your conditions. Graph it and you will be done.

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

Could you tell me how you did it? please?

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

Let us try to see if we can find a parabola that satisfies your conditions: \[ f(x) = a x^2 + b x + c \\ f(0)=0=c\\ c=0\\ f(x) = a x^2 + bx\\ f'(x) = 2 a x + b\\ f'(1)= 2 a + b =0\\ b=-2a\\ f(x)= a x^2 -2 a x\\ f''(x) = 2a < 0\\ \] Take a = -1, or -2 or anything negative will do. So f(x) can be taken \[ f(x)= -x^2 + 2x \]

OpenStudy (anonymous):

So is it the first equation you gave me or the second?

OpenStudy (anonymous):

Or is it the same equation, just rearranged?